OpenStudy (anonymous):

A jewelry box has a volume of x3 + 7x2 + 7x – 15 cm3 and the height is x + 3 cm. Find the polynomial that would represent the area of the bottom of the jewelry box? Explain your reasoning(please).

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

tis a tough 1

OpenStudy (anonymous):

x^3 + 7x^2 + 7x = 15 Or x^3 + 7x^2 + 7x - 15

OpenStudy (anonymous):

would you multiply x^3 + 7x^2 + 7x - 15 by x+3?

OpenStudy (anonymous):

i think you worded it wrong maybe?

OpenStudy (anonymous):

(x+5)*(x-1)*(x+3)

OpenStudy (anonymous):

sorry i was having trouble factoring it :)

OpenStudy (anonymous):

\[ x^3 + 7x^2 + 7x - 15 =x^3+3x^2+4x^2+12x-5x-15\\=x^2(x+3)+4x(x+3)-5(x+3)=(x+3)(x^2+4x-5)\]

OpenStudy (anonymous):

(x+5)*(x-1) is your answer Since x+3 is your height

OpenStudy (anonymous):

Volume=area of bot*height \[(x+3)(x^2+4x−5)=(x+3)*area \ of \ bott\]

OpenStudy (anonymous):

how did u factor it?

OpenStudy (anonymous):

ooh also you could have done algebraic long division using x+3 as denominator.

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

you dont factor you divide use algebraice long division :)

OpenStudy (anonymous):

x^2+4x^2-5x

OpenStudy (anonymous):

If you expand the answer yes :)

OpenStudy (anonymous):

yeyaaaaa ty!

OpenStudy (anonymous):

wait!

OpenStudy (anonymous):

-___-

OpenStudy (anonymous):

x^2+4x-5 :)

OpenStudy (anonymous):

not x^2+4x^2-5x

OpenStudy (anonymous):

oi ty SORRY

OpenStudy (anonymous):

De nada :) see i teaching you spanish also. De nada is you are welcome :D

OpenStudy (anonymous):

haha i know X) gracias mis amigos

OpenStudy (anonymous):

Good job :D

OpenStudy (anonymous):

goodness haha well i must go goodnight! http://www.youtube.com/watch?v=cZO9tMetxno